OpenStudy (barreraa):

Choose the symbol that will create a true sentence when x equals 2. 6 • 8 – 2x _____ 6(5 + x) (Points : 1) < > = ≤

OpenStudy (catlover5925):

> because 6 * 8 = 48 - 4 = 44 now the other side you have to distribute 6 *5 = 30 6 * 2 = 12 so 12 + 30 = 42

OpenStudy (catlover5925):

thanks!!

OpenStudy (barreraa):

so its equal

OpenStudy (anonymous):

Which polynomial identity will prove that 72 = 64 + 8? Difference of Squares Difference of Cubes Sum of Cubes Square of a Binomial can you help me with this @Catlover5925? I think its A

OpenStudy (barreraa):

k thanks

OpenStudy (catlover5925):

@barreraA no your answer is > (greater than)

OpenStudy (catlover5925):

@garciaash i am thinking on yours

OpenStudy (anonymous):

I think it's C not A sorry

OpenStudy (catlover5925):

i think i would agree but im not 100% SURE

OpenStudy (anonymous):

Okay thank you! Could you help me with another one? @Catlover5925

OpenStudy (catlover5925):

yep, open a new thread

OpenStudy (catlover5925):

then tag me